OpenStudy (anonymous):

Given triangle ABC, angle A = 120°, a = 8, b = 3; determine angle B to the nearest whole degree.

OpenStudy (anonymous):

I know the law of cosines is sinA/a = sinB/b = sinC/c

hartnn (hartnn):

do u know sine law ?

hartnn (hartnn):

that was sin law

OpenStudy (anonymous):

This is what I have so far.. sin120/8 = sinB/3 3sin120=8sinB 3/8=sinB I'm lost from there

OpenStudy (anonymous):

Woooow. I meant the law of sines not cosines. I'm just having an off day.

hartnn (hartnn):

use calculator to find arcsin (3/8)

hartnn (hartnn):

wait sin 120 = ... ?

OpenStudy (zarkon):

what is the sine of 120?

OpenStudy (anonymous):

Crap. I'll start over & Thanks so much for helping me

OpenStudy (anonymous):

Okay, now this is what I have... sin120/8=sinB/3 0.58/8=sinB/3 0.07=sinB/3 0.21=sinB sin^-1(0.21) = 12.12 degrees to the nearest whole number is 12. Right?

hartnn (hartnn):

0.58 ? did you use radian mode ? sin 120 = 0.866

OpenStudy (anonymous):

No.. I used degree mode. How do you know when to use each one?

hartnn (hartnn):

120 means its in degrees only. radians angle contains \(\pi\) in it. and sin 120 degrees = 0.866

OpenStudy (anonymous):

Ohh. Well thank you! So my final answer should be 17, right?

hartnn (hartnn):

what u got sin B as ?

OpenStudy (anonymous):

17 degrees

hartnn (hartnn):

why do i get 18.95 :\

OpenStudy (anonymous):

I'm not sure... Lol

hartnn (hartnn):

arcsin 0.324 = 18.95.

OpenStudy (anonymous):

Oh I didn't put the 4 on there so that may have been what messed it up